OpenStudy (anonymous):

Please help! Will fan and give a medal :) Starting from 2.5 miles away, a car drives toward a speed checkpoint and then passes it. The car travels at a constant rate of 66 miles per hour. The distance (d) of the car from the checkpoint is given by d = |2.5 – 66t|. At what times (t) is the car 0.1 miles from the checkpoint? Calculate your answer in seconds.

OpenStudy (anonymous):

Replace d with 0.1 and solve for t.

OpenStudy (anonymous):

\[0.1 =|2.6-66t|\] \[0.1-2.6 = -66t\] \[2.6-0.1 = 66t\] \[\frac{2.6-0.1}{66} = t\]

OpenStudy (anonymous):

Once you solve for t, you will have your answer in hours. You will then need to conver hours to seconds. Do this knowing that there are 3600 seconds in one hour. \[t \times \frac{3600}{1}\] will convert your answer to seconds. Replace t with whatever your answer is.

OpenStudy (anonymous):

Thank you soooo much!

OpenStudy (anonymous):

Is it 5/132 * 3600/1 = 1500/11 ? I'm sorry I am sooo bad at math and still confused lol @walkingdead

OpenStudy (anonymous):

Let's see. Well, first we do 2.6 - 0.1, which is 2.5. Now, we plug into a calculator, 2.5 divided by 66, which is 0.0378 hours. Now, we take that and multiply it by 3600 to convert to seconds, which is 136.37 seconds (approximately).

OpenStudy (anonymous):

Oh, yes, you did it right. :)

OpenStudy (anonymous):

Oh! Oh my gosh thank you! So sorry to bother you again!

OpenStudy (anonymous):

No, no problem! Keep at it and good job!